OpenStudy (anonymous):

How do I write this complex number in a + bi form? √-1 + 2

OpenStudy (anonymous):

2 + i sqrt of - 1 = i

OpenStudy (agreene):

2+i

OpenStudy (anonymous):

What if it was -√-50 - 2 ? Do I factor the 50?

OpenStudy (phi):

yes. \[\sqrt{-50}=\sqrt{-1 \cdot 50}= \sqrt{-1}\cdot \sqrt{50}= i\sqrt{50}\]

OpenStudy (anonymous):

So what would be the written answer as?